OpenStudy (anonymous):

Find the equation of the circle with center at (3, 2) and through the point (5, 4).

OpenStudy (anonymous):

Thank you for helping me

OpenStudy (anonymous):

(y-3)^2+(x-2)^2 = radius^2 radius = distance between (3,2) and (5,4) d = sqrt((y2-y1)^2+(x2-x1)^2) d = sqrt((4-2)^2+(5-3)^2) d = sqrt(2^2+2^2) d =sqrt(4+4) d = sqrt(8) (y-3)^2+(x-2)^2 = (sqrt(8))^2 (y-3)^2+(x-2)^2 = 8